Prove that


i) 𝑑 π‘Žπ‘₯ , π‘Žπ‘¦ = π‘Ž d(x,y)


ii)𝑑 π‘Ž + π‘₯ , π‘Ž + 𝑦 = 𝑑 π‘₯ , 𝑦


where d is a metric induced by on a normed space X

1
Expert's answer
2021-11-30T15:04:03-0500

"d(x,y)=\sum |x_i-y_i|"


i)

"d(ax,ay)=\sum |ax_i-ay_i|=a\sum |x_i-y_i|=ad(x,y)"


ii)

"d(a+x,a+y)=\sum |a+x_i-y_i-a|=\sum |x_i-y_i|=d(x,y)"